On May 1, 2021, Hecala Mining entered into an agreement with the state of New Mexico to obtain the rights to operate a mineraComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 DetComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 CalComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 HowComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 AreComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 Dur
Answer #1

1. The mine will be recorded in the books at it's cost. This cost includes purchase price of rights to use the mine, development cost in preparing the mine and restoration cost to restore the land into it's original condition after mining operation.

Purchase price of the right to use the mine = $ 9,300,000

Development cost in preparing the mine = $ 2,500,000

For finding restoration cost, possible outflows of  restoration cost is given in the problem. The restoration cost will be expected value of this possible outflows.

Therefor restoration cost = (530000*40%)+(630000*30%)+(730000*30%) = $ 620,000

Therefore Cost of mine = $ 9300000+$ 2500000+$ 620000= $ 12,420,000

2. Depreciation and depletion expenses for 2021

units of production method is used for computing depreciation and depletion.

Depreciation or Depletion = Depletion/ Depreciation base * units extracted during the year/ Total units will be extracted

Depletion base for the mine = Cost recorded in the mine = $ 12,420,000

Total tons of ore extracted in the year 2021 = 113000 tons

Total tons of ore extracted in the life of the mine = 730000 tons

therefore Depletion for the mine for the year 2021= 12420000*113000/730000= 1922547.9452

Depreciation base for equipment = Cost of purchase - salvage value at the end of mining process

= $ 143400 -$12000 =$131400

Depreciation for equipment = 131400 *113000/730000 = 20340

Depreciation base for structures = cost of structures for construction = $ 36500

Therefore depreciation for structures for the year 2021 = 36500*113000/730000= 5650

Amount
Depletion $ 1922547.9452
Depreciation expense for equipment $ 20340
Depreciation of structures $ 5650

3.The expected cash outflow on restoration of the mining site is:
530000*40% + 630000*30% + 730000*30% = $620000
The accretion expense to be recorded this period = 620000*1/1.07 4
= $ 472,995.031469

4. The depletion of mine and depreciation of mining equipment and facilities are to
be recorded separately in income statement. Depreciation is provided for plant
assets mainly fixed assets whereas depletion is provided for natural recourses. On
the balance sheet accumulated depreciation appears with the related plant asset
account and accumulated depletion appears with the natural resource asset account.
Depletion is different from depreciation in a way that depletion records the gradual
exhaustion of natural resource reserves while depreciation records wearing out of
tangible assets over its useful life.

5. During 2022 Hecala changed it's estimate in total amount of ore extracted from the mine. So Total units produced should be adjusted . Depletion base also should be changed.

Depreciation or Depletion = Depletion/ Depreciation base * units extracted during the year/ Total units will be extracted

Depletion base for mine = Cost of the mine - depreciation for 2021 = $ 12,420,000-$ 1922547.9452= $ 10497452.0548

Remaining ore can be extracted = New estimation of ore to be extracted - ore extracted during 2021

= 930000- 113000=817000 tons , this will be used in the depletion/depreciation computation for 2022 as total units produced

Ore extracted in 2022= 143000 tons

Therefore Depletion of the mine for 2022 = 10497452.0548*143000/817000= 1837375.3290

Depreciation base for the equipment in 2022 = base in 2021- depreciation in 2021= $131400 -20340=111060

Depreciation of equipment for 2022 = 111060*143000/817000= 19438.8984

Depreciation base for structures for 2022 = cost - depreciation for 2021= $ 36500-5650= $ 30850

Depreciation of structures for 2022 = 30850*143000/817000= 5399.6940

Depreciation for 2022
Depletion $ 1837375.3290
Depreciation expense for equipment $ 19438.8984
Depreciation of structures $ 5399.6940